OpenStudy (anonymous):

3^(x^2) = 9^(5x-8) I get stumped after I get it down to (x^2) = 10x -16

OpenStudy (anonymous):

oh right you did the getting out of the exponent part.

OpenStudy (anonymous):

now solve \[x^2=10x-16\] \[x^2-10x+16=0\] etc

OpenStudy (anonymous):

you did the hard part correctly. now you have a quadratic to solve. this factors even as \[(x-2)(x-8)=0\] and so on

OpenStudy (anonymous):

Thank you!! Well the answer stays in the quadratic right??